# Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation

### v1

@@ Abstract
Direct Preference Optimization (DPO) has been demonstrated to be highly effective in mitigating hallucinations in Large Vision Language Models (LVLMs) by aligning their outputs more closely with human preferences. Despite the recent progress, existing methods suffer from two drawbacks: 1) Lack of scalable token-level rewards; and 2) Neglect of visual-anchored tokens. To this end, we propose a novel Token Preference Optimization model with self-calibrated rewards (dubbed as TPO), which adaptively attends to visual-correlated tokens without fine-grained annotations. Specifically, we introduce a token-level visual-anchored reward as the difference of the logistic distributions of generated tokens conditioned on the raw image and the corrupted one. In addition, to highlight the informative visual-anchored tokens, a visual-aware training objective is proposed to enhance more accurate token-level optimization. Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO. For example, by building on top of LLAVA-1.5-7B, our TPO boosts the performance absolute improvement for hallucination benchmarks.
@@ 1 Introduction
Methods Visual- Anchored Token- level Fine-grained Annotations
@@ 2.1 LVLMs’ Hallucination
Previous studies have mitigated hallucinations by enhancing training data quality, refining decoding strategies, and post-processing generated responses Huang et al. (2024); Leng et al. (2024); Yu et al. (2024a); Han et al. (2024); Chen et al. (2024); Zhou et al. (2023); Yin et al. (2023); Lee et al. (2023); Xia et al. (2024); Shao et al. (2024). While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
@@ 3 Methodology
3.1 Motivation
As illustrated in Figure 1, the generation of each token requires differing levels of integration with visual information. Tokens that are closely correlated with visual inputs are designated as "visual-anchored tokens" in this paper. It is observed that the generation of such visual-anchored tokens can be more susceptible to hallucinations. Furthermore, directly assigning rewards to entire generated sequences can lead to tokens independent of visual information consuming undue optimization resources. This misallocation can reduce the supervisory focus on visual-anchored tokens. To enhance training efficiency and ensure the model attends to relevant visual features, it is crucial to assign token-level rewards based on the reliance on visual information during training.
To achieve these goals, the proposed TPO mitigates hallucination from the following three key aspects: 1) scalable token-level rewards to optimize alignment; 2) automatic identification of visual-anchored tokens by the model itself; 3) self-calibration of reward signals based on visual correlation.
@@ 3.2 Preliminaries
DPO Rafailov et al. (2024) is designed to directly maximize the reward margin between positive and negative responses to align human preferences. Specifically, given a textual input $x$, a visual input $v$, a negative response $y_{l}$, and a preferred positive response $y_{w}$, the reward function is defined as $r(x,v,y_{l}/y_{w})$. According to the Bradley-Terry Bradley and Terry (1952) model, the probability of each response is as follows:
$p(y_{w}>y_{l})=\sigma(r(x,v,y_{w})-r(x,v,y_{l}))$ (1)
where the sigmoid function is denoted as $\sigma(\cdot)$.
In this paper, we focus on the off-policy DPO method, where the probability of preference data is derived from the optimal policy rather than the reward model. The reward function is designed as:
$r(x,v,y)=\beta\log\frac{\pi_{\theta}(y|x,v)}{\pi_{\text{ref}}(y|x,v)}$ (2)
@@
In Equation (2), assuming a response $y\in\{y_{w},y_{l}\}$ has $N$ tokens $(y_{1},...,y_{i},...,y_{N})$, the output distribution of the current policy model $\pi_{\theta}(y|x,v)$ and the reference model $\pi_{\text{ref}}{(y|x,v)}$ can be expressed as $\prod_{y_{i}}p(y_{i}|x,v)$. On this basis, the formulation of a maximum likelihood objective is defined as:
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Bigg{[}\log\sigma$ (3)
$\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}|x,v)}{\pi_{\text{ref}}(y_{w}|x,v)}-\beta\log\frac{\pi_{\theta}(y_{l}|x,v)}{\pi_{\text{ref}}(y_{l}|x,v)}\Bigg{)}\Bigg{]}$
It is demonstrated that DPO assigns equal rewards to all tokens in the responses. When applied to LVLMs for mitigating hallucinations, it lacks emphasis on visual-anchored tokens. Therefore, we propose a novel approach named TPO to mainly solve this problem.
3.3 TPO
3.3.1 Visual-Anchored Rewards
We propose a self-calibrated visual-anchored reward generation mechanism based on Equation (2). On the one hand, it can automatically identify visual-anchored tokens in the response during training. On the other hand, it adaptively and iteratively assigns different rewards for tokens according to their visual reliance. The proposed rewards can not only strengthen the attention to visual information for token generation but also align preferences.
To construct the visual-anchored rewards, we first introduce a score $s$ to measure the visual reliance degree of each token. For this purpose, we add noise into the embedding of the input image $v$ to obtain the corrupted image $v_{c}$:
$v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon$ (4)
Here, we mainly refer Zhou et al. (2024a) and set the noise steps $k=500$, $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$, $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements. [^1] More details can be found in Appendix A. Subsequently, we capture the changes in the probability distribution of each generated token before and after adding noise. For token $y_{i}$ of the response $y$, the calculation of $s_{y_{i}}$ is formulated as Equation (5). Notably, here $p_{log}$ refers to the raw logits output of the model, before applying the softmax transformation.
$s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i})$ (5)
We have shown an example of the distribution of $s$ before and after training in Figure 1, it is shown that visual-anchored tokens will obtain higher scores. In other words, $s$ has the ability to reflect the visual relevance of each token $y_{i}$, this also lays the foundation for the ability of our visual-anchored rewards to self-calibrate based on the visual correlation of tokens.
@@ 3.3.1 Visual-Anchored Rewards
Furthermore, we introduce a self-calibration process to generate the final visual-anchored rewards. This process aims to ensure that positive samples receive higher rewards than negative samples while optimizing the visual relevance of visual-anchored tokens in all responses. Formally, the process is defined as follows, where the sigmoid function $\text{sigmoid}(\cdot)$ is used for normalization.
$c_{y_{i}}=\begin{cases}a+\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{w}\\ a+1-\text{sigmoid}(s_{y_{i}})&\text{if }y_{i}\in y_{l}\end{cases}$ (6)
The calculated $c_{y_{i}}$ is adopted as the final reward in TPO, where $c_{y_{i}}\in(a,a+1)$. In this way, $c$ and $s$ are positively correlated for positive examples, whereas for negative ones, the opposite holds true. In particular, $p(y_{i}|x,v,y_{<i})=p(y_{i}|x,v_{c},y_{<i})$, means that the token does not associate any visual information. In this case, no visual-anchored reward should be assigned for this token. To this end, we set $a=0.5$ in Equation (6), so that when $s=0$, $c=1$, the rewards will not take effect.
3.3.2 Token Preference Optimization
@@
$\max_{\pi}E_{(x,v,y)}\Bigg{[}\,r^{\prime}(x,v,y)-\beta D_{KL}\Bigg{(}\left[\pi_{\theta}(y|x,v)\right.$ (8)
×$\left.\times\pi_{\theta}^{v}(y|x,v)\right]\Bigg{\|}\left[\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)\right]\Bigg{)}\Bigg{]}$
Thus, the optimal solution formula for the maximization objective of the KL-constrained reward is as follows:
$\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)=\frac{1}{Z(x)}\left[\pi_{\text{ref}}(y|x,v)\times\right.$ (9)
$\left.\pi_{\text{ref}}^{v}(y|x,v)\right]\exp\left(\frac{1}{\beta}r^{\prime}(x,v,y)\right)$
Rearranging the Eq (9), we obtain the new reward function:
@@
$=\beta\log\frac{\pi_{\theta}(y|x,v)\times\pi_{\theta}^{v}(y|x,v)}{\pi_{\text{ref}}(y|x,v)\times\pi_{\text{ref}}^{v}(y|x,v)}+\beta Z(x)$
$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log\left(p_{\theta}(y_{i}|x,v,y_{<i})\cdot c_{\theta}\right)\right.$
$\quad\left.-\log\left(p_{\text{ref}}(y_{i}|x,v,y_{<i})\cdot c_{\text{ref}}\right)\right]+\beta Z(x)$
$=\beta\sum_{y_{i}\in\mathcal{Y}}\left[\log p_{\theta}(y_{i}|x,v,y_{<i})-\right.$
$\quad\left.\log p_{\text{ref}}(y_{i}|x,v,y_{<i})+\log\frac{c_{\theta}}{c_{\text{ref}}}\right]+\beta Z(x)$
Compared to the original reward function in DPO (Equation (2)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level. For $c_{\theta}$, its computation is continuously updated at each step during training as the model changes. In the calculation of each token in the entire reward function, we add a term $log\frac{c_{\theta}}{c_{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound. For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease. Due to the different methods of calculating $c_{y_{i}}$ that we set in Equation (6), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
Thus, following the Bradley-Terry model, when given the positive and negative samples $\mathcal{D}=\{x^{(i)},v^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, we obtain our maximum likelihood objective:
$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\bigg{[}\log\sigma$ (11)
$\quad\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}\big{|}x,v)\times\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}(y_{w}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-$
$\quad\beta\log\frac{\pi_{\theta}(y_{l}\big{|}x,v)\times\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}(y_{l}\big{|}x,v)\times\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Bigg{)}\bigg{]}$
@@
Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks. Score and Hall refer to the overall GPT-4 score and hallucination rate, respectively. Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question. The results for POVID Zhou et al. (2024a) and CSR Zhou et al. (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al. (2024) are taken from previous work. (bold: the best score; underline: the second best)
@@ 4.1 Setup
Aligning with previous DPO-based approaches on hallucination mitigation, we adopt the popular LVLM, LLaVA-1.5 Liu et al. (2024b), as the backbone model to validate the effectiveness of our TPO. As for the dataset, we directly utilize the preference pairs provided by RLHF-V (5k) without their fine-grained human annotations.
Benchmarks: We primarily conduct the experiments on three hallucination benchmarks: (1) AMBER Wang et al. (2023): a multi-dimensional hallucination benchmark with over 15k samples. We mainly focus on its discriminative part and report Accuracy and F1 metrics referencing Yu et al. (2024c); (2) MMHal-Bench Sun et al. (2023): it measures the hallucination rate and informativeness of responses. (3) HallusionBench Guan et al. (2024): it evaluates visual illusions and knowledge hallucinations through systematically structured discriminative tasks. We also test the performance on four general evaluation benchmarks: (1) SEED Bench Li et al. (2023a): a benchmark for LVLMs on generative comprehension. (2) MMBench Liu et al. (2025): a comprehensive benchmark designed to evaluate the capabilities across various tasks and modalities. (3) LLaVA Bench Liu et al. (2024c): a benchmark for evaluating multi-modal conversation, detailed description, and complex reasoning. (4) MM-Vet Yu et al. (2023): a benchmark assessing for integrated capabilities.
Baselines: In this work, we mainly compare TPO with the SFT model of LLaVA-1.5-7B, as well as with DPO and V-DPO Xie et al. (2024) methods trained using RLHF-V Yu et al. (2024b) data, along with two improved methods, CSR Zhou et al. (2024b) and POVID Zhou et al. (2024a). Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further compare it with the LLaVA-1.5-13B SFT model, and also against DPO and RLHF-V methods.
4.2 Main Results
[…]
On general benchmarks, compared to other methods, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks. We attribute it to that our method helps the model associate with more visual information when answering questions. This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks.
@@ 4.2 Main Results
Acc F1 Score Hall $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
Table 3: Ablation Studies. Performence of LLaVA-1.5 on hallucination and general benchmarks. (bold: the best score)
Figure 4: Comparison of attention weights for LLaVA-1.5 before and after TPO training. Each horizontal line represents the mean of that data set. The blue section on the left response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g. USB, cloud). The red section on the right answered correctly.

### v3

@@ Abstract
Direct Preference Optimization (DPO) has been demonstrated to be highly effective in mitigating hallucinations in Large Vision Language Models (LVLMs) by aligning their outputs more closely with human preferences. Despite the recent progress, existing methods suffer from two drawbacks: 1) Lack of scalable token-level rewards; and 2) Neglect of visual-anchored tokens. To this end, we propose a novel Token Preference Optimization model with self-calibrated rewards (dubbed as TPO), which adaptively attends to visual-correlated tokens without fine-grained annotations. Specifically, we introduce a token-level visual-anchored reward as the difference of the logistic distributions of generated tokens conditioned on the raw image and the corrupted one. In addition, to highlight the informative visual-anchored tokens, a visual-aware training objective is proposed to enhance more accurate token-level optimization. Extensive experimental results have manifested the state-of-the-art performance of the proposed TPO. For example, by building on top of LLaVA and Qwen, our TPO boosts the performance absolute improvement for hallucination benchmarks.
@@ 1 Introduction
Methods Visual- Anchored Token- level Non Fine-grained Annotations
@@ 2.1 LVLMs’ Hallucination
Previous studies have mitigated hallucinations by enhancing training data quality, refining decoding strategies, and post-processing generated responses Huang et al. (2024); Leng et al. (2024); Yu et al. (2024a); Han et al. (2024); Chen et al. (2024); Zhou et al. (2023); Yin et al. (2023); Lee et al. (2023); Shao et al. (2024). While these methods can lead to more accurate responses, they do not fundamentally resolve the issue of inadequate visual information association in LVLMs.
@@ 3 Methodology
The schematic illustration of the proposed TPO is demonstrated in Figure 2. In Sec. 3.1, we present the preliminaries including the definition and off-policy optimization of DPO. Then we detail the visual-anchored rewards and token preference optimization loss in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.1 Preliminaries
DPO Rafailov et al. (2024) is designed to directly maximize the reward margin between positive and negative responses to align human preferences. Given a textual input $x$, a visual input $v$, a negative response $y_{l}$, and a preferred positive response $y_{w}$, the reward function $r(x,v,y_{l}/y_{w})$ is defined as follows.
$r(x,v,y)=\beta\log\frac{\pi_{\theta}(y|x,v)}{\pi_{\text{ref}}(y|x,v)},$ (1)
@@
where $\pi_{\text{ref}}{(y|x,v)}$ and $\pi_{\theta}(y|x,v)$ respectively represent the reference model and current policy model. On this basis, the formulation of a maximum likelihood objective is defined as:
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (2)
$\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x,v)}{\pi_{\text{ref}}(y_{w}|x,v)}-\beta\log\frac{\pi_{\theta}(y_{l}|x,v)}{\pi_{\text{ref}}(y_{l}|x,v)}\Big{)}\Big{]},$
where $\sigma(\cdot)$ denotes the sigmoid function.
3.2 Visual-Anchored Rewards
Different to the equal confidence for each token in DPO, we propose a visual-anchored by measuring the token-wise visual reliance. Specifically, we firstly add noise into the embedding of the input image $v$ in a total $k$ steps to obtain the corrupted image $v_{c}$:
$v_{c}(k)=\sqrt{\bar{\xi}_{k}}\cdot v+\sqrt{1-\bar{\xi}_{k}}\cdot\epsilon,$ (3)
where $\xi$ is a predefined noise parameter derived from a list with 1,000 equally spaced elements [^1] More details can be found in Appendix A, and experimental analysis can be found in Appendix E. $\bar{\xi}_{k}$ is a cumulant, i.e., $\bar{\xi}_{k}=\prod_{i=0}^{k}\xi_{i}$.
Subsequently, the difference of generated token distribution is computed as follow:
$s_{y_{i}}=p_{log}(y_{i}|x,v,y_{<i})-p_{log}(y_{i}|x,v_{c},y_{<i}),$ (4)
where $s_{y_{i}}$ denotes the distribution difference of the token $y_{i}$ of the response $y$. $p_{log}$ refers to the raw logits output of the model, before applying softmax normalization. One example case is demonstrated in Figure 1, which demonstrates that $s$ reflects the visual relevance of each token $y_{i}$.
Then, a self-calibration process is proposed to generate the final visual-anchored rewards $c_{y_{i}}$.
$c_{y_{i}}=\begin{cases}a+\sigma(s_{y_{i}})&\text{if }y_{i}\in y_{w}\\ a+1-\sigma(s_{y_{i}})&\text{if }y_{i}\in y_{l}\end{cases}$ (5)
@@ 3.2 Visual-Anchored Rewards
where $a$ is a margin value. We set $a=0.5$ in Equation (5), so that when $s=0$, $c=1$, the rewards will not take effect. This process aims to ensure that positive samples receive higher rewards than negative samples while optimizing the visual relevance of visual-anchored tokens in all responses.
3.3 Token Preference Optimization
@@
$\max_{\pi}E_{(x,v,y)}\Big{[}\,r^{\prime}(x,v,y)-\beta D_{KL}\Big{(}\pi_{\theta}(y|x,v)$ (7)
$\cdot\pi_{\theta}^{v}(y|x,v),\pi_{\text{ref}}(y|x,v)\cdot\pi_{\text{ref}}^{v}(y|x,v)\Big{)}\Big{]},$
where $D_{KL}(\cdot,\cdot)$ denotes the KL divergency computation. $\pi_{\theta}^{v}(y|x,v)$ and $\pi_{\text{ref}}^{v}(y|x,v)$ are calculated using the policy model and the reference model, respectively. Thus, the optimal solution formula for the maximization objective of the KL-constrained reward is as follows:
$\pi_{\theta}(y|x,v)\cdot\pi_{\theta}^{v}(y|x,v)=\frac{1}{Z(x,v)}\pi_{\text{ref}}(y|x,v)\cdot$ (8)
$\pi_{\text{ref}}^{v}(y|x,v)\exp\big{(}\frac{1}{\beta}r^{\prime}(x,v,y)\big{)}.$
The partition function of Eq (8) is as follows.
$Z(x,v)=\sum_{y}$ $\pi_{\text{ref}}(y|v,x)\cdot\pi_{\text{ref}}^{v}(y|x,v)$ (9)
$\cdot\exp\big{(}\frac{1}{\beta}r^{\prime}(x,v,y)\big{)}$
Rearranging Eq (8), we obtain the reward function:
@@
$=\beta\log\frac{\pi_{\theta}(y|x,v)\cdot\pi_{\theta}^{v}(y|x,v)}{\pi_{\text{ref}}(y|x,v)\cdot\pi_{\text{ref}}^{v}(y|x,v)}+\beta Z(x,v)$
$=\beta\sum_{y_{i}\in y}\left[\log\left(p_{\theta}(y_{i}|x,v,y_{<i})\cdot{\color[rgb]{1,0,0}c_{y_{i}}^{\theta}}\right)\right.$
$\quad\left.-\log\left(p_{\text{ref}}(y_{i}|x,v,y_{<i})\cdot{\color[rgb]{0,0,1}c_{y_{i}}^{\text{ref}}}\right)\right]+\beta Z(x,v)$
$=\beta\sum_{y_{i}\in y}\Big{[}\log p_{\theta}(y_{i}|x,v,y_{<i})-\log p_{\text{ref}}(y_{i}|x,v,y_{<i})$
$\quad+\log\frac{{\color[rgb]{1,0,0}c_{y_{i}}^{\theta}}}{{\color[rgb]{0,0,1}c_{y_{i}}^{\text{ref}}}}\Big{]}+\beta Z(x,v),$
where $c_{y_{i}}^{\theta}$ and $c_{y_{i}}^{\text{ref}}$ represent the token reward calculated using the policy model and the reference model, respectively.
Compared to the original reward function in DPO (Eq (1)), we multiply each $p(y_{i}|x,v,y_{<i})$ by the generated visual-anchored rewards $c_{y_{i}}$ at the token level. $c_{y_{i}}^{\theta}$ is continuously updated at each step during training as the model changes. To calculate each token in the entire reward function, we add a term $log\frac{c_{y_{i}}^{\theta}}{c_{y_{i}}^{\text{ref}}}\in(-log3,log3)$, which has a reasonable upper and lower bound. For positive samples, this term is expected to increase, while for negative samples, it is expected to decrease. Due to the different methods of calculating $c_{y_{i}}$ that we set in Eq (5), this will encourage the increase of $s_{y_{i}}$ during the training process, making the token generation focus more on visual information.
Thus, following the Bradley-Terry model, when given the positive and negative samples $\mathcal{D}=\{x^{(k)},v^{(k)},y_{w}^{(k)},y_{l}^{(k)}\}_{k=1}^{N}$, we obtain our maximum likelihood objective:
$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (11)
$\quad\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}\big{|}x,v)\cdot\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}(y_{w}\big{|}x,v)\cdot\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-$
$\quad\beta\log\frac{\pi_{\theta}(y_{l}\big{|}x,v)\cdot\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}(y_{l}\big{|}x,v)\cdot\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Big{)}\Big{]}$
$=\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})+\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$
$\quad\Big{(}\beta\log\frac{\pi_{\theta}^{v}(y_{w}\big{|}x,v)}{\pi_{\text{ref}}^{v}(y_{w}\big{|}x,v)}-\beta\log\frac{\pi_{\theta}^{v}(y_{l}\big{|}x,v)}{\pi_{\text{ref}}^{v}(y_{l}\big{|}x,v)}\Big{)}\Big{]}$
According to Eq (10), we can deduce as follows.
$\mathcal{L}_{TPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim D}\Big{[}\log\sigma$ (12)
$\quad\Big{(}\beta\sum_{y_{w_{i}}\in y_{w}}\Big{[}\log\left(p_{\theta}(y_{w_{i}}|x,v,y_{w_{<i}})\right)$
$\quad-\log p_{\text{ref}}(y_{w_{i}}|x,v,y_{w_{<i}})+\log\frac{c_{y_{w_{i}}}^{\theta}}{c_{y_{w_{i}}}^{\text{ref}}}\Big{]}$
$+\sum_{y_{l_{i}}\in y_{l}}\Big{[}\log\left(p_{\theta}(y_{l_{i}}|x,v,y_{l_{<i}})\right)$
$\quad-\log p_{\text{ref}}(y_{l_{i}}|x,v,y_{l_{<i}})+\log\frac{c_{y_{l_{i}}}^{\theta}}{c_{y_{l_{i}}}^{\text{ref}}}\Big{]}\Big{)}\Big{]}$
where $c_{y_{w_{i}}}^{\theta}$ and $c_{y_{w_{i}}}^{\text{ref}}$ represent the token reward calculated for $y_{w}$ using the policy model and the reference model, respectively. The same applies to $c_{y_{w_{i}}}^{\theta}$, $c_{y_{w_{i}}}^{\text{ref}}$ and $y_{l}$.
@@
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
Table 2: Performence of LLaVA-1.5 on hallucination and general benchmarks. Score and Hall refer to the overall GPT-4 Achiam et al. (2023) score and hallucination rate, respectively. Easy represents the accuracy of with original images, hard represents the accuracy with manually edited challenging images, and aAcc is the average accuracy for each question. The results for POVID Zhou et al. (2024a) and CSR Zhou et al. (2024b) are based on our testing of their open-source model weights, while the results for V-DPO Xie et al. (2024) are taken from previous work (bold: the best score; underline: the second best)
Method AMBER MMHal HallusionBench
[…]
Qwen2-vl 86.5 90.0 3.5 29.0 67.0 48.8 64.0
+DPO 86.5 90.0 3.7 28.1 67.3 49.3 64.5
+TPO 86.4 89.9 4.2 18.8 67.9 50.0 65.2
Table 3: Performence of Qwen2-vl-7B on hallucination benchmarks.
@@ 4.1 Setup
Aligning with previous DPO-based approaches on hallucination mitigation, we mainly adopt the popular LVLM, LLaVA-1.5 Liu et al. (2024b), as the backbone model to validate the effectiveness of our TPO. Furthermore, to evaluate the effectiveness of TPO on more advanced and powerful model, we implement TPO training based on Qwen2-vl Wang et al. (2024), and compare it with the DPO method. For the dataset, we directly utilize the preference pairs provided by RLHF-V (5K) without their fine-grained human annotations.
Benchmarks
We primarily conduct the experiments on three hallucination benchmarks: AMBER Wang et al. (2023), MMHal-Bench Sun et al. (2023), and HallusionBench Guan et al. (2024). In this section, we mainly focus on AMBER’s discriminative task and report the accuracy and F1 metrics referencing Yu et al. (2024c). In addition, we provide the results of its Chair metric in Appendix D. Moreover, we also evaluate the performance of TPO on four general benchmarks: SEED Bench Li et al. (2023a), MMBench Liu et al. (2025), LLaVA Bench Liu et al. (2024c) and MM-Vet Yu et al. (2023). These benchmarks are used to evaluate the performance of the models on general tasks after hallucination alignment.
Baselines
We mainly compare TPO with the LLaVA-1.5-7B SFT model, as well as with the DPO and V-DPO Xie et al. (2024) methods trained using RLHF-V Yu et al. (2024b) data, along with two improved methods, CSR Zhou et al. (2024b) and POVID Zhou et al. (2024a). Moreover, to evaluate the effectiveness and robustness of TPO as the model size increases, we further evaluate the performance of TPO on the LLaVA-1.5-13B model and compared it with DPO. Additionally, to demonstrate the advantages of TPO, we reproduced the strong baseline method, RLHF-V, on LLaVA-1.5-13B and conducted a comparison.
To further demonstrate the effectiveness of TPO on more advanced and powerful model, we additionally employ Qwen2-vl-7B as the baseline mode and compare our TPO with DPO.
@@
Acc F1 Score Hal $\downarrow$ Easy Hard aAcc SEED MMB LLaVA MM-Vet
@@
Table 4: Ablation Studies. Performence of LLaVA-1.5 on hallucination and general benchmarks.
Average score Noun/Adj Others
Ground Truth 1.83 0.90
Ground Truth (TPO) 5.72 4.87
Response of LLaVA 1.48 0.83
Response of LLaVA+TPO (TPO) 5.67 4.59
Table 5: Average score from Equation 4 of Noun/Adj tokens and other tokens. Here, Ground Truth and Ground Truth (TPO) represent the scores calculated for the ground truth answer using LLaVA-1.5-7B and LLaVA-1.5-7B+TPO. Response of LLaVA and LLaVA+TPO (TPO) correspond to the outputs before and after TPO training and the scores calculated by LLaVA-1.5-7B and LLaVA-1.5-7B+TPO, corresponding to Fiure 1.
4.2 Main Results
[…]
On general benchmarks, our approach remains stable against the original LLaVA model and achieves the greatest improvement on most benchmarks. We attribute it to that our method helps the model associate with more visual information when answering questions. This shows that our approach can improve hallucination issues while maintaining good performance on general evaluation tasks. The results also indicate that TPO introduces performance fluctuations on the LLaVABench and MMVet metrics for LLaVA-1.5-13B. This occurs because the training data primarily focuses on aligning preferences within the context of hallucinations, which inevitably leads to slight overfitting. Although augmenting the dataset with task-specific data could mitigate this issue, it is not the research focus of this paper.
4.3 Results on Qwen2-vl
As Table 3 shown, we report the results on the key metrics of three hallucination benchmarks.
The results indicate that our TPO outperforms DPO on most benchmarks. On Qwen2-vl, which has strong inherent capabilities, using 5K RLHF-V data for DPO alignment barely improves the performance. However, introducing TPO leads to a significant further enhancement. This not only demonstrates the effectiveness of TPO on other powerful models but also shows that TPO’s token-level visual anchor rewards brings higher data utilization efficiency. Although Qwen2-vl shows almost no room for improvement on AMBER, TPO still achieves significant improvement on the Chair metric, as shown on Figure 5. TPO can capture and learn more subtle preferences from the data.
Figure 3: Comparison of attention weights for LLaVA before and after TPO training. Each horizontal line represents the mean of that data. The blue section response incorrectly, with many ’visual-anchored tokens’ tokens having high attention weights but resulting in hallucinated responses (e.g. USB). The red section on the right answered correctly.